Please help me solve this equation:
The length of a rectangular room is 3 feet less than twice its width. Find the dimensions of the room if the perimeter is 54 feet.

Answer by checkley77(12844)

LENGTH=2W+3
WIDTH=W
2L+2W=PERIMETER
2(2W+3)+2W=54
4W+6+2W=54
6W=54-6
6W=48
W=48/6
W=8 IS THE WIDTH.
L=2*8+3=16=3=19 IS THE LENGTH.
PROOF
2*19=2*8=54
38+16=54
54=54